# Learning Dynamic Bayesian Networks from Data: Foundations, First Principles and Numerical Comparisons

### v1

@@ 3.4 Sample Complexity for Forecasting
Where the intuition described above arises is in recent results in sample complexity. We shall see that while the arithmetic of (13) is still correct for DBNs, there are indeed important distinctions on the sample complexity with respect to the number of different trajectories $N$ and the length of the trajectory $T$.
Classically, theoretical analyses of time series sample complexity typically assumed that the trajectory is much longer than the mixing time and by cutting the synthetic burn in period, as such obviates any need to analyze historical dependence. (see the review of the previous results in [64])
We shall report on the theoretical small sample complexity results reported in [64], which is yet unpublished but extends and otherwise mentions similar recent results in [68, 67, 73, 16].
@@
We report on the one prominent result as far as learning so as to achieve accurate inference on BNs. The classic work [15] reports on a sample complexity, in VC dimension analysis, of modeling a Bayesian Network to be,
@@ 3.5 Sample Complexity for Identification
In general, identifying the Bayesian Network is NP-Complete with respect to the number of variables [10]. It is noted that the number of possible DAGs for 10 variables is greater than 4$\times$1018 [52].
@@
The work [49] presents poly-time identifiability in the case of bounded treewidth or acyclic super-structure, and otherwise confirms NP-Hardness of search with respect to data. A creative recent work [25] uses models from physiology to argue for $O\left(M^{k}\right)$ practical complexity, with $M$ the cardinality of a discrete valued network and $k$ is the number of potential parents.
@@
Finally, the work [66] considers sample complexity of causal discovery specifically, which runs at the number of samples required being $O\left(n!l^{3n/8}\right)$, where $l$ is the cardinality of the possible random variable values in a discrete network.
@@ 3.6 Formalization: Open System Forecasting Model
For instance, DBNs are often used for predictive maintenance, as in [1, 69]. By an appropriate representation of the underlying complex engineering system as distilled into some low dimensional latent structure, one can develop DBNs to monitor signals of deterioration or damage in the system as based on the historical transitions over time in performance.
An interesting formalism of this is given in [21]. For some underlying stochastic process $\dot{Y}=g(Y(t),W(t))$ with (e.g., Brownian) noise $W(t)$ where $Y\in\mathcal{Y}$ is very high, if not infinite, dimensional, one can consider a DBN model as a finite dimensional reduced order model of the system, and one that maximizes the information relevance towards maintenance. Formal guarantees are provided as far as probabilistic invariance, that is,
@@
indicating the potential for DBNs to serve as useful indicators of higher level properties of stochastic processes, regardless of the fundamental impossibility of formal causal structure identification in such cases.
4 Understanding Structure and Parameter Learning Algorithms
Here we present a number of important points of consideration that can be observed from looking at the literature at successful attempts at representation as far as inference and learning. With the distinctions described below, we are able to properly identify and characterize existing structure-weight learners, as well as suggest and provide straightforward extensions to fill in the natural empty places in the taxonomy.
Structure Learning
Parameter Learning
[…]
Frequentist and Bayesian
We shall use the frequentist versus Bayesian distinction to indicate a point estimate based on the optimization of a loss function or criterion, and a probabilistic model, implemented with sampling, that obtains a posterior distribution of the structure and weights given the data, respectively. A frequentist estimate is given as a complete specific structure encoding and a specific value for the parameters. It is generally expected, or at least sought, that the relationships that the graph identifies between various is statistically significant. This presumption often becomes unrealistic in practice, and obtaining an appropriately scaled statistically significant entire network, that is, with all significant edges, is typically unavailable. However, since DBNs are generative rather than discriminative, this is often not a practical concern, as they are a component in an overall statistical modeling pipeline.
An effective and commonly used technique is to employ mixtures of a finite set of structures, see [24]. This provides flexibility and the transparent uncertainty in the model, without having to mix through the entire combinatorial space defining possible structure.
For instance, consider a particular point estimate of a structure and set of parameters. However, consider that the set of parameters is close to zero, and moreover, that is so close so as to include zero in a, e.g., 95% confidence interval. In this case, it is clear that this implies that the presence of this edge itself in the graph is suspect, that is, not implied by the data.
Hierarchical and One-Shot Methods
In general one can consider most learning methods to be hierarchical in the sense of first learning the structure, and with an amortized structure estimating or sampling the weights. The use of SEMs defined by adjacency matrices including both structure and weights simultaneously introduced what can be referred to as a one-shot approach (we remark the interestingly similar recent popularity of one-shot methods for neural architecture search as including parameter learning [29]).
In this case, a point estimate is obtained for both the structure and the weights simultaneously by solving an appropriate optimization problem that fits both of these as decision variables to the data. To this end there are two approaches we see in the literature. In [45] an IP (for BNs, readily adapted to DBNs) is presented that treats the structure as binary variables encoding the activation of edge links in the graph and the parameters as separate variables, and solves the challenging nonlinear mixed IP (relaxation into conic programs was considered in [37]). Alternatively, the recent work DYNOTEARS [2] presented a gradient based method for solving the structure-parameter learning as a purely continuous optimization problem for weight matrices in the graph. Enforcing sparsity is done to encourage proper structure learning.
We make one additional remark going back to hierarchical approaches. Note that one can consider that the frequentist-Bayesian distinction can be applied to present a taxonomy of methods. As a curious example, many Bayesian scoring methods, e.g. the IP method [3], can be considered hybrid frequentist-Bayesian. This is because implicitly the grading is done with a Bayesian parameter model, but a point estimate, that is one unique structure, is returned. Methodologically, we see that the advantages of a hierarchical is an offline calculation of scoring that permits the use of simple and powerful off the shelf commercial grade IP solvers, and the disadvantage is the conceptual contradiction of applying a frequentist mindset to learning structure with Bayesian models as weights. However, one can easily mitigate this in practice by sampling from multiple structures, as weighted in frequency by their respective marginal likelihoods. Regardless, theoretically, in the asymptotic regime, consistency can still be maintained with all approaches and variations thereof, however [36].
@@ 5.1 Criteria
In order to ascertain the performance of different structures, a score function serves as an objective in an optimization process. The score function is meant to evaluate the statistical accuracy of a model. In performing structure learning as guided by a score, we are performing a likelihood, or some maximum a posteriori maximization, in the process of traversing the decision landscape of structures.
Selection criteria for models appears in both the BN/PGM and the time series modeling literature. In [47] a thorough exploration of the evaluation and computation of various criteria is presentented for a range of different time series models. In [19] it is recommended to use a general form for an information criterion to evaluate possible networks is, for $N$ samples and $k$ parameters:
@@
There are a variety of options as far as how to use this criterion to choose a model. We first present a few that are natural but do not appear consistently in the literature, before continuing to discuss the Bayesian structure learning approach.
We write generically the likelihood $L_{k}(\{X\}|\Theta,\mathcal{G})$, and write,
  1. 1.
    One Shot Frequentist Directly maximize $L_{k}$ with respect to $\Theta$ and $\Xi$ simultaneously for a simultaneous frequentist solution, with $C_{k,N}$ defined as a sparsity metric (i.e., $l0$ “norm”)
  2. 2.
    Hierarchical Frequentist Maximize $L_{k}(\{X\}|\Theta(\mathcal{G}),\mathcal{G})$ with respect to $\mathcal{G}$. To evaluate the likelihood given $\mathcal{G}$, one must compute $\Theta(\mathcal{G})$. This itself can be the maximum likelihood of the parameters, i.e.,
@@
    where conditioning on $\mathcal{G}$ enforces certain components of $\theta$ to be zero.
A popular alternative is to use Bayesian criteria. In this case, the actual score function is the marginal posterior of the candidate structure given the data, that is $p(\mathcal{G}|\{X\})$. For particular kinds of parametrized DBNs, computing this posterior can be done in closed form. For a classic discussion on the statistical intuition, motivation, and some formulations of Bayesian criteria, see [31]
@@ 5.2 Likelihood Calculations
A common assumption made in the literature  [36, 26] is that of global parameter independence. That is, it holds that the parameters $(\theta|\bar{\mathcal{G}})$ can be decomposed to be separable across the transitions for each variable $X_{i}$, i.e. using the notation $\theta^{\bar{\mathcal{G}},i}$ to indicate parameters associated with the transition step for variable $X_{i}(t+1)$,
@@
Notice that here the separability is with respect to trajectories, and not necessarily time steps. Furthermore, below we shall see that an additional assumption of local independence is needed to furthermore assure independence across the parameters defining the dependence of the transition of $X_{i}$ on each parent.
Since the likelihood is a separable function of the parameters, maximizing it corresponds to maximizing the set of parameters separately for each , that is, seek to maximize, where we perform the usual condition dependence chain $p(X(t+1),X(t),...,X(0)|\theta)=p(X(t+1)|X(t),X(t-1),...,X(1)|\theta)=p(X(t+1)|X(t),\theta)p(X(t)|X(t-1),\theta),...,p(X(0))$ to facilitate the presentation of the chain of conditioning to facilitate the posterior derivation.
@@ 5.2.1 Binary Variables
Indeed this corresponds to the local parameter independence and the unrestricted multinomial conditions that facilitates the closed form computation for the Bayesian Dirichlet scores. To this end, we now extend the presentation in [31] (see also [30]) to include the contribution of the static $Z$ variables to the model.
@@
Now we present the computation of the Bayesian Dirichlet scores. This amounts to computing the marginal posterior of the structure by performing an integration treating parameter as nuisance. This is derived, for instance, in [26], and used in the popular integer BN structure learner GOBNILP [14]. The marginal posterior of the structure is given by:
@@ 5.2.2 Gaussian DBNs
Now we present the derivation of the likelihood and Bayesian criterion (BGe) for DBNs with Gaussian models. The development follows [26] and extends their derivation in two ways. First we perform the recursion for computing the entire trajectory time data. Second, we include a specific parametrization and show how one can simultaneously perform the recursion to obtain a posterior of the weights. We, however, simplify our model to only include Markovian influence, and not lagged autoregressive effects.
We apply the model in [26] to (10) to obtain the following transition likelihood function for the first step and prior for both the overall likelihood transition and the parameters themselves:
@@
$\begin{array}[]{l}p\left(\mu(0)|W,\bar{\mathcal{G}}\right)=\mathcal{N}\left(\nu(0),\alpha_{\mu}W\right)\\ \nu(0)=\begin{pmatrix}\mu^{+}(0):=\eta_{0}+\eta_{d}\cdot\mu^{x}_{j\in dpa_{d}}(0)+\eta_{s}\cdot\mu^{x}_{j\in dpa_{s}}+\eta_{z}\cdot\mu^{z}_{j\in dpa_{z}(i)}(0)\\ \mu^{x}_{j\in dpa_{d}}(0)\\ \mu^{x}_{j\in dpa_{s}}(0)\\ \mu^{z}_{j\in dpa_{z}(i)}(0)\end{pmatrix}\\ p\left(W|\bar{\mathcal{G}}\right)=c(n_{i}(1),\alpha_{w})|T|^{\alpha_{w}/2}|W|^{(a_{\alpha}-n_{i}(1)-1)/2}e^{-1/2\mathop{tr}(TW)}\equiv\text{Wishart}\left(W|\alpha_{w},T\right)\\ c(n_{i}(1),\alpha_{w}):=\left(2^{\alpha_{w}n/2}\pi^{n(n-1)/4}\prod\limits_{i=1}^{n_{x}}\Gamma\left(\frac{\alpha_{w}+1-i}{2}\right)\right)^{-1}\\ n(1)=1+|dpa(i)|\\ \alpha_{\mu}W=\\ \psi\begin{pmatrix}(\Upsilon(0)\,/\,\Upsilon_{\setminus 0}(0))(\Upsilon(0)\,/\,\Upsilon_{\setminus 0}(0))^{T}&\mu_{dpa_{d}(i)}^{x}(0)(\Upsilon(0)\,/\,\Upsilon_{\setminus 0}(0))(\Upsilon(0)\,/\,\Upsilon_{\setminus d}(0))^{T}&..\\ ..&..&..\\ ..&..&\mu_{dpa_{z}(i)}^{z}(0)(\Upsilon(0)\,/\,\Upsilon_{\setminus 0}(0))(\Upsilon(0)\,/\,\Upsilon_{\setminus d}(0))\end{pmatrix}\end{array}$ (29)
where $A/B$ denotes the Schur complement of $A$ with respect to $B$. In [26, Theorem 4 and Theorem 5] it is shown that parameter independence is preserved through the computation of the posterior. Note that the posterior now is with respect to all the data that is present in a transition. The DAG structure ensures that $\nu(0)$ is well defined as a vector, rather than implicitly as a function of $\mu^{x}_{j\in dpa_{s}(i)}(0)$. Finally, the last line related the two models together, indicating how the Wishart distribution arises from the parameter distribution, in this case.
@@
Now, with this redundant embedding in both prior in the variable space and parameter space, we deduce how to compute the posterior of the distribution distribution of the data $\mu$ and $W$ from [26] for $T=2$, and subsequently, compute the posterior of the parameters in the model, while showing it is equivalent by a straightforward Bayesian posterior propagation. After deriving the base case $T=2$, we continue with the induction for $T$ to $T+1$, in order to derive the final posterior of the data, from which we can compute the marginal likelihood of the structure, as well as sample the final posterior values.
@@
In general terms, the form is broadly preserved. As such, we can reproduce the evaluation for the marginal likelihood, that is the BGe score, directly from [26]
@@
This defines the distribution of weights. Let us finally consider the reverse transformation, of obtaining the score from the weights. Indeed, this can be done straightforwardly, as then $W(T)$ can be recovered from $\Upsilon(T)$, and then $\mu(T)$ will have mean $\nu(T)$

### v2

@@ 3.5 Sample Complexity for Forecasting
Where the intuition described above arises is in recent results in sample complexity. We shall see that while the arithmetic of (14) is still correct for DBNs, there are indeed important distinctions on the sample complexity with respect to the number of different trajectories $N$ and the length of the trajectory $T$ in certain data regimes.
Classically, theoretical analyses of time series sample complexity typically assumed that the trajectory is much longer than the mixing time and by cutting the synthetic burn in period, as such obviates any need to analyze historical dependence. (see the review of the previous results in [tu2022learning])
We shall report on the theoretical small sample complexity results reported in [tu2022learning], which is yet unpublished but extends and otherwise mentions similar recent results in [xing2022identification, xin2022learning, zheng2020non, dean2020sample].
@@
We report on the one prominent result as far as learning so as to achieve accurate inference on BNs. The classic work [dasgupta1997sample] reports on a sample complexity, in VC dimension analysis, of modeling a Bayesian Network to be,
@@ 3.6 Sample Complexity for Identification
In general, identifying the Bayesian Network is NP-Complete with respect to the number of variables [chickering1996learning]. It is noted that the number of possible DAGs for 10 variables is greater than 4$\times$1018 [peters2017elements].
@@
The work [ordyniak2013parameterized] presents poly-time identifiability in the case of bounded treewidth or acyclic super-structure, and otherwise confirms NP-Hardness of search with respect to data. A creative recent work [geduk2022practical] uses models from physiology to argue for $O\left(M^{k}\right)$ practical complexity, with $M$ the cardinality of a discrete valued network and $k$ is the number of potential parents.
@@
Finally, the work [wadhwa2021sample] considers sample complexity of causal discovery specifically, which runs at the number of samples required being $O\left(n!l^{3n/8}\right)$, where $l$ is the cardinality of the possible random variable values in a discrete network.
@@ 3.7 Formalization: Open System Forecasting Model
For instance, DBNs are often used for predictive maintenance, as in [amin2019fault, yu2013novel]. By an appropriate representation of the underlying complex engineering system as distilled into some low dimensional latent structure, one can develop DBNs to monitor signals of deterioration or damage in the system as based on the historical transitions over time in performance.
An interesting formalism of this is given in [esmaeil2015dynamic]. For some underlying stochastic process $\dot{Y}=g(Y(t),W(t))$ with (e.g., Brownian) noise $W(t)$ where $Y\in\mathcal{Y}$ is very high, if not infinite, dimensional, one can consider a DBN model as a finite dimensional reduced order model of the system, and one that maximizes the information relevance towards maintenance. Formal guarantees are provided as far as probabilistic invariance, that is,
@@
indicating the potential for DBNs to serve as useful indicators of higher level properties of stochastic processes, regardless of the fundamental impossibility of formal causal structure identification in such cases. These considerations, and the study of DBNs without causal sufficiency in general, is broadly limited, however.
@@ 4.1 Criteria
In order to ascertain the goodness of fit of different DBN graph structures, a score function serves as an objective in an optimization process. The score function is meant to evaluate the statistical accuracy of a model together with its parsimony. In performing structure learning as guided by a score, we are performing a likelihood, or some maximum a posteriori maximization, in the process of traversing the decision landscape of structures.
Selection criteria for models appears in both the BN/PGM as well as the time series modeling literature. In [mcquarrie1998regression] a thorough exploration of the evaluation and computation of various criteria is presented for a range of different time series models. In [douc2014nonlinear] it is recommended to use a general form for an information criterion to evaluate possible networks is, for $N$ samples (not distinguishing between trajectories and time points) and a total $k$ parameters:
@@
From an alternative perspective, we consider how the Likelihood can be computed and maximized. Writing the quantity as $L_{k}(\{X\}|\Theta,\mathcal{G})$ we see that it depends on both the structure $\mathcal{G}$ and the parameters $\Theta$. Observing that 1) the parameter space itself depends on the graph structure, and 2) the likelihood is differentiable with respect to the parameters whereas the structure is a discrete quantity that presents, fundamentally, a combinatorial optimization problem.
A natural approach, which can be described as a sort of hierarchical frequentist method, would be to maximize $L_{k}(\{X\}|\Theta(\mathcal{G}),\mathcal{G})$ with respect to $\mathcal{G}$. In doing so, however, to evaluate the likelihood given $\mathcal{G}$, one must compute $\Theta(\mathcal{G})$ that is,
@@
where conditioning on $\mathcal{G}$ enforces the appropriate structure of $\theta$ to be zero. Computationally, one can also pursue a one shot approach wherein the structure and weights are maximized simultaneously, with some appropriate functional enforcement of consistency of the structure with the zero sparsity structure of weights. While formally, a global maximum obtained with a hierarchical procedure is going to be the same equivalent pair $(\mathcal{G}^{*},\Theta^{*})$, in practice the choice does carry a distinction as far as tendencies and patterns of the solutions of the optimization procedure.
From a frequentist perspective, the term $C_{k,N}$ is a regularizer chosen instrumentally to prioritize sparser solutions. Sparse basis recovery has long been an important area of statistical research (e.g. [wainwright2009information], with a popular monograph (with a chapter on PGMs) [wainwright2019high] (see also [wainwright2008graphical]).
In taking a Bayesian perspective, one can consider a functional / computational equivalence by considering $C_{k,N}$ as the prior, with varying strength corresponding to the degree of confidence in the prior. Formally, the score function is the marginal posterior of the candidate structure given the data, that is $p(\mathcal{G}|\{X\})$. This, of course, is unknown in form but can be appropriately decomposed with the parameters serving as “nuisance” vectors, i.e.,
[…]
For particular kinds of parametrized DBNs, computing this posterior can be done in closed form without evaluating explicitly the terms involving $\theta$. For a classic discussion on the statistical intuition, motivation, and some formulations of Bayesian criteria, see [heckerman1995learning].
@@ 4.2 Likelihood Calculations
A common assumption made in the literature  [koller2009probabilistic, geiger2002parameter] is that of global parameter independence. That is, it holds that the parameters $(\theta|\bar{\mathcal{G}})$ can be decomposed to be separable across the transitions for each variable $X_{i}$, i.e. using the notation $\theta^{\bar{\mathcal{G}},i}$ to indicate parameters associated with the transition step for variable $X_{i}(t+1)$,
@@
This assumption significantly eases computation, as the solution can be obtained by maximizing the likelihood for the set of parameters separately for each $i$. Now, to obtain the expression for the total likelihood in this case, perform the usual condition dependence chain
$\begin{array}[]{l}p(X(t+1),X(t),...,X(0)|\theta)=p(X(t+1)|X(t),X(t-1),...,X(1)|\theta)\\ \qquad=p(X(t+1)|X(t),\theta)p(X(t)|X(t-1),\theta),...,p(X(0))\end{array}$
This yields the complete expression below, where we now include all the structured dependencies.
@@
We will also use $\bar{\Xi}_{i}$ to indicate the variables corresponding to the set of dependencies $dpa(i)$
@@ 4.2.1 Binary Variables
Indeed this corresponds to the local parameter independence and the unrestricted multinomial conditions that facilitates the closed form computation for the Bayesian Dirichlet scores. To this end, we now extend the presentation in [heckerman1995learning] (see also [heckerman2008tutorial]) to include the contribution of the static $Z$ variables to the model.
@@
Now we present the computation of the Bayesian Dirichlet scores. This amounts to computing the marginal posterior of the structure by performing an integration treating parameter as nuisance. This is derived, for instance, in [geiger2002parameter], and used in the popular integer BN structure learner GOBNILP [cussens2020gobnilp]. The marginal posterior of the structure is given by:
@@ 4.2.2 Gaussian DBNs
Now we present the derivation of the likelihood and Bayesian criterion (BGe) for DBNs with Gaussian models. The development follows [geiger2002parameter] and extends their derivation in two ways. First we perform the recursion for computing the entire trajectory time data. Second, we include a specific parametrization and show how one can simultaneously perform the recursion to obtain a posterior of the weights. We, however, simplify our model to only include Markovian influence, and not lagged autoregressive effects.
We apply the model in [geiger2002parameter] to (11) to obtain the following transition likelihood function for the first step and prior for both the overall likelihood transition and the parameters themselves:
@@
$\begin{array}[]{l}p\left(\mu(0)|W,\bar{\mathcal{G}}\right)=\mathcal{N}\left(\nu(0),\alpha_{\mu}W\right)\\ \nu(0)=\begin{pmatrix}\mu^{+}(0):=\eta_{0}+\eta_{d}\cdot\mu^{x}_{j\in dpa_{d}}(0)+\eta_{s}\cdot\mu^{x}_{j\in dpa_{s}}+\eta_{z}\cdot\mu^{z}_{j\in dpa_{z}(i)}(0)\\ \mu^{x}_{j\in dpa_{d}}(0)\\ \mu^{x}_{j\in dpa_{s}}(0)\\ \mu^{z}_{j\in dpa_{z}(i)}(0)\end{pmatrix}\\ p\left(W|\bar{\mathcal{G}}\right)=c(n_{i}(1),\alpha_{w})|T|^{\alpha_{w}/2}|W|^{(a_{\alpha}-n_{i}(1)-1)/2}e^{-1/2\mathop{tr}(TW)}\equiv\text{Wishart}\left(W|\alpha_{w},T\right)\\ c(n_{i}(1),\alpha_{w}):=\left(2^{\alpha_{w}n/2}\pi^{n(n-1)/4}\prod\limits_{i=1}^{n_{x}}\Gamma\left(\frac{\alpha_{w}+1-i}{2}\right)\right)^{-1}\\ n(1)=1+|dpa(i)|\\ \alpha_{\mu}W=\\ \psi\begin{pmatrix}U(0)&U_{d}(0)&..\\ ..&..&..\\ ..&..&U_{z}(0)\end{pmatrix}\\ U(0)=(\Upsilon(0)\,/\,\Upsilon_{\setminus 0}(0))(\Upsilon(0)\,/\,\Upsilon_{\setminus 0}(0))^{T}\\ U_{d}(0)=\mu_{dpa_{d}(i)}^{x}(0)(\Upsilon(0)\,/\,\Upsilon_{\setminus 0}(0))(\Upsilon(0)\,/\,\Upsilon_{\setminus d}(0))^{T}\\ U_{z}(0)=\mu_{dpa_{z}(i)}^{z}(0)(\Upsilon(0)\,/\,\Upsilon_{\setminus 0}(0))(\Upsilon(0)\,/\,\Upsilon_{\setminus d}(0))\\ ...\end{array}$ (30)
where $A/B$ denotes the Schur complement of $A$ with respect to $B$. In [geiger2002parameter, Theorem 4 and Theorem 5] it is shown that parameter independence is preserved through the computation of the posterior. Note that the posterior now is with respect to all the data that is present in a transition. The DAG structure ensures that $\nu(0)$ is well defined as a vector, rather than implicitly as a function of $\mu^{x}_{j\in dpa_{s}(i)}(0)$. Finally, the last line related the two models together, indicating how the Wishart distribution arises from the parameter distribution, in this case.
@@
Now, with this redundant embedding in both prior in the variable space and parameter space, we deduce how to compute the posterior of the distribution distribution of the data $\mu$ and $W$ from [geiger2002parameter] for $T=2$, and subsequently, compute the posterior of the parameters in the model, while showing it is equivalent by a straightforward Bayesian posterior propagation. After deriving the base case $T=2$, we continue with the induction for $T$ to $T+1$, in order to derive the final posterior of the data, from which we can compute the marginal likelihood of the structure, as well as sample the final posterior values.
@@
In general terms, the form is broadly preserved. As such, we can reproduce the evaluation for the marginal likelihood, that is the BGe score, directly from [geiger2002parameter]
@@
This defines the distribution of weights. Note that obtaining the score from the weights can be done straightforwardly as well, as then $W(T)$ can be recovered from $\Upsilon(T)$, and then $\mu(T)$ will have mean $\nu(T)$